A block of jello with mass 8.08 kg is resting on an inclined plane. The plane is inclined to the horizontal by an angle 28.3°. The jello has a coefficient of static friction with respect to the surface of the plane of µs=0.77. What is the magnitude of the static friction force between the jello and the ramp?
Answer #1

magnitude of static friction = us*N

in an Inclined plane N = mg cos

Ff = uS * mg cos

Ff = 0.77*8.08*9.8*cos 28.3

= 53.68 N

so the magnitude of static friction is 53.68 N